OpenStudy (pandora):

I got this far and can't finish it: find the limit as x goes to zero of. Sin pi/6 (cos delta x) + cos pi/6 (sin delta x) - 1/2. All over delta x

OpenStudy (across):

Is this what you mean?\[\lim_{\Delta x\to0}\frac{\sin(\pi/6)\cos(\Delta x)+\cos(\pi/6)\sin(\Delta x)}{\Delta x}\]

OpenStudy (pandora):

Yeah sorry I'm new to this and dont know how to do that notation

OpenStudy (amistre64):

you got that far, from where did you find that?

OpenStudy (amistre64):

split the fraction into 2 parts\[\frac{a+b}{n}=\frac an+\frac bn\]

OpenStudy (pandora):

This is the original problem. Lim of delta x goes to zero of. Sin (pi/6 + delta x ) + 1. All over Delta x

OpenStudy (amistre64):

then you used the angle addition property and .. yeah

OpenStudy (pandora):

Uh u lost me there

OpenStudy (amistre64):

i was just catching up to the part you already did; but still, you just need to split the fraction in two

OpenStudy (pandora):

Oh gotcha but then once the fractions are split what do you do. Can you divide by the delta x

OpenStudy (amistre64):

you recall the definitions for:\[\lim_{h\to\ 0}\frac{sin~h}{h}\text{ and }\lim_{h\to\ 0}\frac{cos~h}{h}\]

OpenStudy (amistre64):

the sin goes to 1; and the cos goes to 0

OpenStudy (pandora):

Oh duh yeah the lightbulb just went on. Thanks so much for being patient wi me

OpenStudy (amistre64):

:) good luck